OpenStudy (xximhisgirlxx):

A line segment has one endpoint at P(2, 5) and another, Q, on the line x = 3. Which expression represents the slope of line segment PQ? y-5 x-3 2x+5y 3x+5y @jim_thompson5910

jimthompson5910 (jim_thompson5910):

point Q is the point (3,y) since it is on the line x = 3

jimthompson5910 (jim_thompson5910):

now use the slope formula to get m = (y2-y1)/(x2-x1) m = (y-5)/(3-2) m = (y-5)/(1) m = y-5 So the slope is y-5

OpenStudy (xximhisgirlxx):

thank you

jimthompson5910 (jim_thompson5910):

np